OpenStudy (anonymous):

What is the vertex of the parabola x = 5(y – 4)2 – 1 ?

OpenStudy (anonymous):

(–1, 4) @jim_thompson5910 am i right ?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

the only math you need is you+me=forever

OpenStudy (anonymous):

Lol . and thankyou @0cx42

OpenStudy (anonymous):

could you help with a few more ?

OpenStudy (anonymous):

yes i could just baby put your lips together and come real close and blow my whistel baby

OpenStudy (anonymous):

What is the center and radius of the circle with the equation x2 + y2 – 10x – 12y + 36 = 0 ? Center: (5, 6); Radius = 4 or Center: (5, 6); Radius = 5

OpenStudy (anonymous):

Center (5,-6) radius = 7

OpenStudy (anonymous):

Wait

OpenStudy (anonymous):

ok. i thhought the first one , but im not sure.

OpenStudy (anonymous):

It's the second one

OpenStudy (anonymous):

Center (5,6) Radius=5

OpenStudy (anonymous):

thqankyou(: . do you know how to graph ?

OpenStudy (anonymous):

Yes X=5 Y=6

OpenStudy (anonymous):

noo lol im not graphing this question . its another question.

OpenStudy (anonymous):

Oh ok

OpenStudy (anonymous):

before that. do you know how to do this last one ? what are the vertices of the hyperbola given by the equation (y-8)^2/9-(x-5)^2/4=1

OpenStudy (anonymous):

You will have Hyperbola's postive and negative

OpenStudy (anonymous):

One is up one is down

OpenStudy (anonymous):

I got 5 and 11

OpenStudy (anonymous):

and 5,5 right ?

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

ok now graphing . i just have a few. graph \[\frac{x ² }{ 49 } + \frac{ y² }{ 9 }\]

OpenStudy (anonymous):

i dont know how to use a graphing calculator either.

OpenStudy (anonymous):

helloo ?